# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ II.3 Exact solutions of zero modes at smooth domain walls
with exponents $\alpha_{L}^{\pm}=-s{{v}}_{L}\pm{q}_{L}$, $\alpha_{R}^{\pm}=s{{v}}_{R}\pm{q}_{R}$, where ${q}_{L,R}=\sqrt{{v}_{L,R}^{2}+{m}_{L,R}}$, which depend only on the values of the fields at large distances $|x|\gg w$, and where the functions $F_{1,2}$ are bounded or diverge polynomially for $x\to\pm\infty$. The detailed derivation of these solutions is given in Appendix A. Asymptotically, these solutions give
@@
$\varphi_{1,2}^{s}(x\to+\infty)$ $\sim{e}^{-\alpha_{R}^{\mp}|x|}={e}^{-\mu_{R}^{\mp}|x|}{e}^{-\mathrm{i}\kappa_{R}^{\mp}|x|},$ (8b)
where $\mu_{L,R}^{\pm}=1/\xi_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ are the decay rates and decay lengths, while $\kappa_{L,R}^{\pm}=2\pi/\lambda_{L,R}^{\pm}=\mathrm{Im}(\alpha_{L,R}^{\pm})$ the momentum and wavelengths describing the oscillatory behavior of the solutions. Obviously $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s\mathrm{Re}({v}_{L,R})$ if $\mathrm{Re}({q}_{L,R})=0$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})$ if $\mathrm{Im}({v}_{L,R})=0$. If the system becomes asymptotically hermitian for $x\to\mp\infty$, i.e., $\mathrm{Im}({m}_{L,R})=\mathrm{Im}({v}_{L,R})=0$ for $x\to\mp\infty$, then $\kappa_{L,R}^{+}=\kappa_{L,R}^{-}=0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}\geq 0$, with $\mu_{L,R}^{+}=\mu_{L,R}^{-}$ for ${{v}}_{L,R}^{2}+{m}_{L,R}=0$, while $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s{v}_{L,R}$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})\neq 0$ for ${{v}}_{L,R}^{2}+{m}_{L,R}<0$.
Let us assume now that the fields ${v}(x)$ and ${m}(x)$ can be expanded in powers of $y(x)$ up to the first and second order, respectively, i.e., as ${v}(x)={v}_{0}+{v}_{1}y(x)$, and ${m}(x)={m}_{0}+{m}_{1}y(x)+{m}_{2}y(x)^{2}$, where the coefficients of the expansion are given by $v_{0}=v_{L}$, ${v}_{1}={v}_{R}-{v}_{L}$, and $m_{0}=m_{L}$, $m_{1}+m_{2}={m}_{R}-{m}_{L}$, ${m}_{2}=2({m}_{L}+{m}_{R})-4{m}(0)$. In this case, we found that the general solution of the modified Jackiw-Rebbi equation can be written in terms of hypergeometric functions. Indeed, for a given pseudospin $s=\pm 1$, we found that
@@
where ${{\,}_{2}F_{1}}(a,b,c,y)$ is the hypergeometric function with
$a_{1,2}$ =±$=\pm w\left({{q}_{L}}-{{q}_{R}}\right)+\tfrac{1}{2}+\tfrac{1}{2}{\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}}}$ (10a)
$b_{1,2}$ =±$=\pm w\left({{q}_{L}}-{{q}_{R}}\right)+\tfrac{1}{2}-\tfrac{1}{2}{\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}}}$ (10b)
@@
assuming $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$ (we denote $\mathbb{Z}_{\leq}$ the set of nonpositive integers). The detailed derivation of this solution is given in Appendix B. These solutions generalize to the nonhermitian case the solutions derived in Ref. marra_topological_2024, . The asymptotic behavior of the functions $F_{1,2}$ is summarized in Table 1. If either $c_{1}-a_{1}=c_{2}-a_{2}\in\mathbb{Z}_{\leq}$ or $c_{1}-b_{1}=c_{2}-b_{2}\in\mathbb{Z}_{\leq}$, the solutions of the hypergeometric equation become more complicated, since some of the hypergeometric functions become linearly dependent or ill-behaved for $x\to\pm\infty$. We will not address these cases explicitly in this work.
@@
The solutions $\varphi_{1,2}^{s}(x)$ simplify in the cases where $|{v}_{L,R}|={v}$ or ${m}_{L,R}={m}$ or if the fields are uniform, as shown in Table I of Ref. marra_topological_2024, . Since the functions $F_{1,2}$ are bounded or diverge polynomially, as summarized in Table 1, the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$ uniquely determine the asymptotic behavior of the solutions. Hence, the existence of zero modes, i.e., particular solutions that are normalizable and satisfy the boundary conditions on a given interval, is uniquely and solely determined by the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$. Since these exponents only depend on $v_{L,R}$, $m_{L,R}$, one can infer that the values of the fields at large distances and the boundary conditions univocally determine the existence and number of zero modes and their pseudospin $s$ on infinite and semi-infinite intervals.
@@ III.1 Topological invariant
In hermitian systems, the bulk-boundary correspondence hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 establishes a relation between the number of topologically protected modes at the boundary between two topologically inequivalent phases. For uniform and hermitian fields $m(x)=m\in\mathbb{R}$, $v(x)=v\in\mathbb{R}$, one can define the topological invariant $W\in\mathbb{Z}_{2}$ characterizing the Hamiltonian as $W=\operatorname{sgn}(v)$ for $m\leq 0$ and $W=0$ otherwise kitaev_unpaired_2001 . Hence, assuming uniform fields at the asymptotes $x\to\pm\infty$, the phase on the left $x\ll-w$ has topological invariant $W_{L}=\operatorname{sgn}(v_{L})$ if $m_{L}\leq 0$ and $v_{L}\neq 0$, and $W_{L}=0$ otherwise. Analogously, the phase on the right $x\gg w$ gives $W_{R}=\operatorname{sgn}(v_{R})$ if $m_{R}\leq 0$ and $v_{R}\neq 0$, and $W_{R}=0$ otherwise (see Refs. kitaev_unpaired_2001 ; shen_topological_2011 ). The number of topologically protected zero-energy modes localized at the boundary is equal to the difference between the values of the topological invariant on the right and on the left of the boundary $\Delta W=|W_{L}-W_{R}|$.
Here, we want to generalize the bulk-boundary correspondence to nonhermitian systems, particularly to nonhermitian systems described by nonhermitian modified Jackiw-Rebbi equations as in Eq. 1. These nonhermitian systems can be seen as an adiabatical "deformation" of the conventional hermitian systems. Hence, the bulk gap of the hermitian Hamiltonian is transformed into a line gap in the nonhermitian case schindler_hermitian_2023 . We will define here a topological invariant for these nonhermitian Hamiltonians by an adiabatic deformation of the topological invariant in the hermitian case, justified by our analysis of the decay rates of the localized boundary modes solutions of the Jackiw-Rebbi equations in Eq. 1.
To determine the existence, number, and asymptotic properties of the zero modes, in agreement with our previous considerations on the line gap nonhermitian topology and the definition of bulk topological invariant in Eq. 6, we define
@@
where we use the common convention $\operatorname{sgn}(0)=0$. We will refer to $W_{L,R}$ and $M_{L,R}$ as the (nonhermitian) topological invariant and the (nonhermitian) topological mass on the left and right side of the smooth domain wall, respectively. These quantities depend only on the values of the fields ${m}_{L,R}$ and ${v}_{L,R}$ at large distances $|x|>w$. For the asymptotically hermitian case $\mathrm{Im}({m}_{L,R})=\mathrm{Im}({v}_{L,R})=0$, one has that $\operatorname{sgn}({M}_{L,R})=\operatorname{sgn}({m}_{L,R})$, ${K}_{L,R}=|\mathrm{Im}({q}_{L,R})|$, and ${W}_{L,R}=\operatorname{sgn}({v}_{L,R})$ for ${m}_{L,R}\leq 0$ and ${W}_{L,R}=0$ otherwise. Hence, our definitions of nonhermitian topological mass $M_{L,R}$ and nonhermitian topological invariant $W_{L,R}$ in Eqs. 11 and 11c reduces to the usual definition of topological mass and topological invariant in the hermitian case.
@@
$\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=\left\{\begin{array}[]{ll}\pm 1,&\text{ for }{M}_{L}>0,\\ 0\text{ or }-sW_{L},&\text{ for }{M}_{L}=0,\\ -sW_{L},&\text{ for }{M}_{L}<0.\end{array}\right.$ (12)
In the first case, the solution $\varphi_{1}^{s}(x)$ decays while the solution $\varphi_{2}^{s}(x)$ diverges for $x\to-\infty$. In the second case, one of the solutions $\varphi_{1,2}^{s}(x)$ converges to a constant value. In the last case, both solutions $\varphi_{1,2}^{s}(x)$ decays for $x\to-\infty$ as long as $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$. Analogous statements hold for $\alpha_{R}^{\pm}$ by exchanging left and right ${v}_{L}\to{v}_{R}$, ${m}_{L}\to{m}_{R}$, $W_{L}\to W_{R}$, ${M}_{L}\to{M}_{R}$, and $s\to-s$.
We are now ready to demonstrate the nonhermitian generalization of the bulk-boundary correspondence. To do so, we will demonstrate that the number of localized zero-energy modes in the interval $-\infty<x<\infty$ (infinite system) equals $|W_{L}-W_{R}|$, and that the number of localized zero-energy modes in the interval $0\leq x<\infty$ (semi-infinite system) is equal to $|W_{R}|$.
III.2 Infinite interval $-\infty<x<\infty$
Zero modes on the interval $-\infty<x<\infty$ satisfying the boundary conditions $\varphi(\pm\infty)=0$ are given by the modes $\varphi_{1,2}^{s}(x)$ in Eq. 7 with $\mathrm{Re}(\alpha_{L}^{\pm})>0$ and $\mathrm{Re}(\alpha_{R}^{\pm})>0$. In this case, there exists $|W_{L}-W_{R}|$ zero modes, i.e., there are two independent modes if $W_{L}W_{R}=-1$, one independent mode if $W_{L}=\pm 1$ and $W_{R}=0$ or if $W_{R}=\pm 1$ and $W_{L}=0$. To demonstrate these statements, we need to find the particular solutions that satisfy $\varphi(x\to\pm\infty)=0$, i.e., $\varphi(y\to 0)=\varphi(y\to 1)=0$. Assuming $|F_{1,2}(y)|$ convergent or logarithmically divergent at $y=0,1$, we need to care about the factors $y^{w\alpha_{L}^{-}}$ and $(1-y)^{w\alpha_{R}^{-}}$, which converge to zero at $y=0$ and $y=1$ only if $\mathrm{Re}({\alpha_{L}^{-}})>0$ and $\mathrm{Re}({\alpha_{R}^{-}})>0$, respectively. This gives only four possible outcomes, discussed below and summarized in Table 2.

### v2

@@ II.3 Exact wavefunctions
for a given pseudospin $s=\pm 1$, and with $\alpha_{L}^{\pm}=-s{{v}}_{L}\pm{q}_{L}$, $\alpha_{R}^{\pm}=s{{v}}_{R}\pm{q}_{R}$, where ${q}_{L,R}=\sqrt{{v}_{L,R}^{2}+{m}_{L,R}}$. A fundamental property of the exponents $\alpha_{L,R}^{\pm}$ is that they only depend on the asymptotic values of the scalar fields $|x|\gg w$. The functions $F_{1,2}$ are bounded or diverge polynomially for $x\to\pm\infty$. The detailed derivation of these solutions is given in Appendix˜A. The asymptotic behavior of these eigenmodes is
@@
$\varphi_{1,2}^{s}(x\to+\infty)$ $\sim{e}^{-\alpha_{R}^{\mp}|x|}={e}^{-\mu_{R}^{\mp}|x|}{e}^{-\mathrm{i}\kappa_{R}^{\mp}|x|}.$ (7b)
Here, $\mu_{L,R}^{\pm}=\mathrm{Re}(\alpha_{L,R}^{\pm})$ are the decay rates on the opposite sides of the smooth domain wall, and $\xi_{L,R}^{\pm}=1/\mathrm{Re}(\alpha_{L,R}^{\pm})$ the corresponding decay lengths; $\kappa_{L,R}^{\pm}=\mathrm{Im}(\alpha_{L,R}^{\pm})$ are the momenta describing the oscillatory behavior of the eigenmodes on the opposite sides of the smooth domain wall, and $\lambda_{L,R}^{\pm}=2\pi/\mathrm{Im}(\alpha_{L,R}^{\pm})$ the corresponding wavelengths. Obviously $\mu_{L,R}^{+}=\mu_{L,R}^{-}=\mp s\mathrm{Re}({v}_{L,R})$ if $\mathrm{Re}({q}_{L,R})=0$ and $\kappa_{L,R}^{+}=-\kappa_{L,R}^{-}=\mathrm{Im}({q}_{L,R})$ if $\mathrm{Im}({v}_{L,R})=0$.
We can further assume that the scalar fields can be expanded in powers of $y(x)$. Specifically, we assume ${v}(x)={v}_{0}+{v}_{1}y(x)$ (up to the first order) and ${m}(x)={m}_{0}+{m}_{1}y(x)+{m}_{2}y(x)^{2}$ (up to the second order). Here, the coefficients of the expansion can be related to the values of the fields at large distances as $v_{0}=v_{L}$, ${v}_{1}={v}_{R}-{v}_{L}$, and $m_{0}=m_{L}$, $m_{1}+m_{2}={m}_{R}-{m}_{L}$, ${m}_{2}=2({m}_{L}+{m}_{R})-4{m}_{D}$ and to the value of the mass at the center of the domain wall ${m}_{D}={m}(0)$. With these further assumptions, we get that the eigenmodes are given by Eq.˜6 with
@@
where ${{\,}_{2}F_{1}}(a,b,c,y)$ is the hypergeometric function and
$a_{1,2}$ $=\pm w\left({{q}_{L}}-{{q}_{R}}\right)+\tfrac{1}{2}+\tfrac{1}{2}{\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}}}$ (9a)
$b_{1,2}$ $=\pm w\left({{q}_{L}}-{{q}_{R}}\right)+\tfrac{1}{2}-\tfrac{1}{2}{\sqrt{\left(2ws{v}_{1}-1\right)^{2}+4w^{2}{m}_{2}}}$ (9b)
@@
assuming $c_{1,2}-a_{1,2},c_{1,2}-b_{1,2}\notin\mathbb{Z}_{\leq}$ (we denote by $\mathbb{Z}_{\leq}$ the set of nonpositive integers). The detailed derivation of this solution is given in Appendix˜B. These solutions generalize to the nonhermitian case the solutions derived in Ref. marra_topological_2024 . The asymptotic behavior of the functions $F_{1,2}$ is summarized in Table˜1. If either $c_{1}-a_{1}=c_{2}-a_{2}\in\mathbb{Z}_{\leq}$ or $c_{1}-b_{1}=c_{2}-b_{2}\in\mathbb{Z}_{\leq}$, the solutions of the hypergeometric equation become more complicated, since some of the hypergeometric functions become linearly dependent or ill-behaved for $x\to\pm\infty$. We will not address these cases explicitly in this work.
$F_{1,2}(x\to\mp\infty)=1$
$a_{1,2}\in\mathbb{Z}_{\leq}$ or $b_{1,2}\in\mathbb{Z}_{\leq}$ $F_{1,2}(x\to\pm\infty)=\frac{\Gamma(c_{1,2})\Gamma(c_{1,2}-a_{1,2}-b_{1,2})}{\Gamma(c_{1,2}-b_{1,2})\Gamma(c_{1,2}-a_{1,2})}\neq 0$
$\mathrm{Re}({q}_{R,L})>0$, $a_{1,2},b_{1,2}\notin\mathbb{Z}_{\leq}$ $F_{1,2}(x\to\pm\infty)=\frac{\Gamma(c_{1,2})\Gamma(2w{q}_{R,L})}{\Gamma(c_{1,2}-a_{1,2})\Gamma(c_{1,2}-b_{1,2})}\neq 0$
$\mathrm{Re}({q}_{R,L})=0$, ${q}_{R,L}\neq 0$, $a_{1,2},b_{1,2}\notin\mathbb{Z}_{\leq}$ $F_{1,2}(x\to\pm\infty)\sim{e}^{\mp\mathrm{i}\mathrm{Im}(2{q}_{R,L})x}\frac{\Gamma(c_{1,2})\Gamma(-2w{q}_{R,L})}{\Gamma(a_{1,2})\Gamma(b_{1,2})}+\frac{\Gamma(c_{1,2})\Gamma(2w{q}_{R,L})}{\Gamma(c_{1,2}-a_{1,2})\Gamma(c_{1,2}-b_{1,2})}$
${q}_{R,L}=0$, $a_{1,2},b_{1,2}\notin\mathbb{Z}_{\leq}$ $F_{1,2}(x\to\pm\infty)\sim\pm(x/w)\frac{\Gamma(c_{1,2})}{\Gamma(a_{1,2})\Gamma(b_{1,2})}$
@@
The eigenmodes in the case where one or both of the scalar fields are uniform or for some special symmetric cases such as when $|{v}_{L,R}|={v}$ or ${m}_{L,R}={m}$ can be found in Table I of Ref. marra_topological_2024 . Since the functions $F_{1,2}$ are bounded or diverge polynomially, as summarized in Table˜1, the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$ uniquely determine the asymptotic behavior of the eigenmodes. Hence, the existence of zero modes, i.e., particular solutions that are normalizable and satisfy the boundary conditions on a given interval, is uniquely and solely determined by the exponents $\alpha_{L}^{\pm}$ and $\alpha_{R}^{\pm}$. Since these exponents only depend on $v_{L,R}$, $m_{L,R}$, one can infer that the existence and number of zero energy modes and their pseudospin $s$ is univocally determined by the values assumed by the scalar fields at large distances.
@@ III.1 Topological invariant
As we have seen, the general solutions of the Jackiw-Rebbi equation in the nonhermitian regime are a direct generalization of the solutions in the hermitian regime obtained in our previous work marra_zero_2025 . However, determining the existence and number of zero modes in the nonhermitian regime, i.e., determining the existence and number of particular solutions satisfying the boundary conditions, is not as straightforward.
In the hermitian regime, the bulk-boundary correspondence hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 determines a relation between topologically protected modes at the boundary between two topologically inequivalent phases. For uniform and hermitian fields $m(x)=m\in\mathbb{R}$, $v(x)=v\in\mathbb{R}$, the topological invariant $W\in\mathbb{Z}_{2}$ is $W=\operatorname{sgn}(v)$ (nontrivial) for $m\leq 0$ and $W=0$ (trivial) otherwise kitaev_unpaired_2001 . Hence, assuming uniform fields at the asymptotes $x\to\pm\infty$, the phase on the left $x\ll-w$ has topological invariant $W_{L}=\operatorname{sgn}(v_{L})$ if $m_{L}\leq 0$ and $v_{L}\neq 0$, and $W_{L}=0$ otherwise. On the other hand, the phase on the right $x\gg w$ gives $W_{R}=\operatorname{sgn}(v_{R})$ if $m_{R}\leq 0$ and $v_{R}\neq 0$, and $W_{R}=0$ otherwise (see Refs. kitaev_unpaired_2001 ; shen_topological_2011 ). The difference between the topological invariants on the opposite sides of the boundary $\Delta W=|W_{L}-W_{R}|$ gives the number of topologically protected zero-energy modes localized at the boundary.
Here, we will generalize the bulk-boundary correspondence to nonhermitian systems, particularly to nonhermitian systems described by nonhermitian regime of the modified Jackiw-Rebbi equation. These nonhermitian systems can be seen as an adiabatical "deformation" of the conventional hermitian systems. Hence, the bulk gap of the hermitian Hamiltonian is transformed into a line gap in the nonhermitian case schindler_hermitian_2023 . We will define here a topological invariant for these nonhermitian Hamiltonians by an adiabatic deformation of the topological invariant in the hermitian case, justified by our analysis of the decay rates of the localized boundary modes of the modified Jackiw-Rebbi equation in Eq.˜1.
In agreement with our previous considerations on the line gap nonhermitian topology and the definition of bulk topological invariant in Eq.˜5, we define
@@
where we use the common convention $\operatorname{sgn}(0)=0$. We will refer to $W_{L,R}$ and $M_{L,R}$ as the (nonhermitian) topological invariant and the (nonhermitian) topological mass on opposite sides of the smooth domain wall. These quantities depend only on the asymtpotic values of the scalar fields ${m}_{L,R}$ and ${v}_{L,R}$. For the asymptotically hermitian case $\mathrm{Im}({m}_{L,R})=\mathrm{Im}({v}_{L,R})=0$, one has that $\operatorname{sgn}({M}_{L,R})=\operatorname{sgn}({m}_{L,R})$, ${K}_{L,R}=|\mathrm{Im}({q}_{L,R})|$, and ${W}_{L,R}=\operatorname{sgn}({v}_{L,R})$ for ${m}_{L,R}\leq 0$ and ${W}_{L,R}=0$ otherwise. Hence, our definitions of nonhermitian topological mass $M_{L,R}$ and nonhermitian topological invariant $W_{L,R}$ in Eq.˜10 reduces to the usual definition of topological mass and topological invariant in the hermitian case.
@@
[…]
In the first case, the solution $\varphi_{1}^{s}(x)$ decays while the solution $\varphi_{2}^{s}(x)$ diverges for $x\to$-$\infty$. In the second case, one of the solutions $\varphi_{1,2}^{s}(x)$ converges to a constant value. In the last case, both solutions $\varphi_{1,2}^{s}(x)$ decays for $x\to$-$\infty$ as long as $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$. Analogous statements hold for $\alpha_{R}^{\pm}$ by exchanging left and right ${v}_{L}\to{v}_{R}$, ${m}_{L}\to{m}_{R}$, $W_{L}\to W_{R}$, ${M}_{L}\to{M}_{R}$, and $s\to-s$, that is
[…]
We are now ready to demonstrate the nonhermitian generalization of the bulk-boundary correspondence. To do so, we will demonstrate that the number of localized zero-energy modes in an infinite-size system defined on the interval $(-\infty,\infty)$ is equal to $|W_{L}-W_{R}|$, and that the number of localized zero-energy modes in a semi-infinite system defined on the interval $0,\infty)$ is equal to $|W_{R}|$.
III.2 Infinite interval $(-\infty,\infty)$
We now consider zero-energy modes in an infinite-size system. These are defined as the eigenmodes with zero energy that satisfy the boundary conditions $\varphi(\pm\infty)=0$ on the interval $(-\infty,\infty)$. In order to satisfy the boundary conditions, the modes need to decay asymptotically for $x\to\pm\infty$, which mandates that the exponents need to be $\mathrm{Re}(\alpha_{L}^{\pm})>0$ and $\mathrm{Re}(\alpha_{R}^{\pm})>0$. In this case, there exists $|W_{L}-W_{R}|$ zero modes, i.e., two modes if $W_{L}W_{R}=-1$, and one mode if $|W_{L}|=1$ and $W_{R}=0$ or $|W_{R}|=1$ and $W_{L}=0$. To demonstrate these statements, we need to find the particular solutions that satisfy $\varphi(x\to\pm\infty)=0$, i.e., $\varphi(y\to 0)=\varphi(y\to 1)=0$. Assuming $|F_{1,2}(y)|$ convergent or logarithmically divergent at $y=0,1$, we need to care about the factors $y^{w\alpha_{L}^{-}}$ and $(1-y)^{w\alpha_{R}^{-}}$, which converge to zero at $y=0$ and $y=1$ only if $\mathrm{Re}({\alpha_{L}^{-}})>0$ and $\mathrm{Re}({\alpha_{R}^{-}})>0$, respectively. This gives only four possible outcomes, discussed below and summarized in Table˜2.